OpenStudy (anonymous):

what does 3 sqrt(u) (sqrt(448)+4sqrt(7) equal or factor to? I got 24 Usqrt(7)

OpenStudy (anonymous):

\(448=2^6\times 7\) should help

OpenStudy (zehanz):

Do you mean:\[3 \sqrt{u}(\sqrt{448}+4\sqrt{7})\] I can see that \[\sqrt{448}=\sqrt{64 \cdot 7}=8\sqrt{7}\] Add up to see how may times sqrt(7) you've got...

OpenStudy (anonymous):

\(8\sqrt{7}+4\sqrt{7}=12\sqrt{7}\)

OpenStudy (zehanz):

It's more than 24 times...

OpenStudy (anonymous):

wait what is more than 24 times

OpenStudy (zehanz):

Well, you've got:\[3\sqrt{u} \cdot 12 \sqrt{7}\] That is \[36\sqrt{7u}\]

OpenStudy (anonymous):

o i see thank you